OpenStudy (anonymous):

\[sin^{-1}0\] Let's see.... \[sin{-1}x=y\] \[x=siny\] \[o=siny\] where from here?

OpenStudy (anonymous):

ask yourself this "do i know a number (angle) between \(-\frac{\pi}{2}\) and \(\frac{\pi}{2}\) whose sine is zero? "

OpenStudy (anonymous):

you do not need all that stuff above. like if i ask what is the square root of 25, you say 5 because you know that \(5^2=25\)

OpenStudy (anonymous):

if you see \(\sin^{-1}(0)\) you know that is \(0\) because you know \(\sin(0)=0\)

OpenStudy (anonymous):

Let's see here I know that sin of zero is zero. The same holds for the inverse?

OpenStudy (anonymous):

and if you see \(\sin(\frac{\sqrt{2}}{2})\) you know it is \(\frac{\pi}{4}\) because \[\sin(\frac{\pi}{4})=\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

the inverse just bothers me I guess

OpenStudy (anonymous):

on the graph of \(\sin(x)\) is \((0,0)\) so if you switch the ordered pairs you still get \((0,0)\)

OpenStudy (anonymous):

on the graph of \(\sin(x)\) is the point \((\frac{\pi}{2},\frac{\sqrt{2}}{2})\) so on the graph of the inverse is the point \((\frac{\sqrt{2}}{2},\frac{\pi}{4})\)

OpenStudy (anonymous):

Yes! Yep that makes sense. So when we do the inverse we do (a,b) -> (b,a) Makes sense

OpenStudy (anonymous):

right i made a typo above, ignore it

OpenStudy (anonymous):

yeah I noticed. No worries. Thanks though @satellite73

OpenStudy (anonymous):

do koala's receive medals too?